" The larger of two numbers is 20 more than the smaller. Four times the larger is 70 more than 5 times the smaller. Find the numbers."
Answer by rapaljer(4551)

Let x = smaller
x+20 = larger

"Four times the larger is 70 more than 5 times the smaller. Find the numbers."
4(x+20) = 5*x + 70
4x+80 = 5x + 70

Subtract 4x from each side:
80=x+70

Subtract 70 from each side:
10=x

x=10 Smaller
x+20 = Larger
10+20= 30 Larger

Check:
"Four times the larger is 70 more than 5 times the smaller. Find the numbers."
4*30 = 5*10 + 70
120=50+70

It checks.

R^2